OpenStudy (anonymous):

3a-(-1)=-5 How to solve?

OpenStudy (anonymous):

minus a negative is the same as adding. So, 3a + 1 = -5

OpenStudy (anonymous):

then, move the 1 to the right. then, divide the whole equation by 3.

OpenStudy (anonymous):

3a = -5-1 3a = -6 a = -2

OpenStudy (anonymous):

thanks! i got it